# Phase diagram of the quantum spin-1/2 Heisenberg-$\Gamma$ model on a frustrated zigzag chain

### v1

@@ III.4 Phase diagram
Figure 11: (a) Local magnetization $\langle S^{z}_{i}\rangle$ (upper panel) in the 1/3-plateau phase, clearly showing the up-up-down spin configuration. The peak wave number $q$ of ${\cal S}^{zz}(q)$ is shown as a function of $h$, where, within $0.5\pi\leq k\leq\pi$ we find $q=2\pi/3$. (a) The spin correlation function, $\langle S^{x}_{i}S^{x}_{i+N/2}\rangle\neq 0$ (i=25), as a function of $h$ that takes finite value in the fluctuating-F phase at $1.5\lesssim h\lesssim 2.15$.
@@ IV Effect of magnetic field
It is known that the $J_{1}$-$J_{2}$ Heisenberg model in the AF-AF and F-AF regions exhibits a rich phase diagram in an applied magnetic fieldHikihara et al. (2008, 2010). Here, our $\Gamma$-terms change these phase diagrams a lot. We consider an external magnetic field along the $z$-axis whose Hamiltonian is given as ${\cal H}+{\cal H}_{f}$ using Eq.(1) with
@@ IV.1 Magnetic phase diagram
Figure 9 shows the magnetic phase diagram at fixed values of $\theta=0.05\pi,0.1\pi,0.3\pi,0.5\pi$. We first focus on small $\Gamma$, namely $\theta=0.05\pi,0.1\pi$. Compared to the zigzag Heisenberg model, the $\Gamma$-term transforms some of the nonmagnetic phases to magnetic ordered ones. For example, the fluctuating-F phase at around $(\phi,\theta,h)\sim(0.2,0.05,1.8)$ in Fig. 9(a) was originally a TLL2 phaseHikihara et al. (2010) but was stabilized as an emergent small magnetic order in the $xy$ plane. Furthermore, in the F-AF region $0.5\pi<\phi<\pi$, the canted UUDD phase where up-up-down-down magnetic order emerges in the $xy$ plane, was originally a vector chiral, nematic, and other multipolar phases when $\Gamma=0$.
At large $\Gamma$, i.e. $\theta=0.3\pi,0.5\pi$ in Figs. 9(c) and 9(d), the nonmagnetic phases like TLL and ND are almost wiped out. Magnetic phases, F and UD, that appeared in the $h=0$ phase diagram are preserved while the moments are canted off the $xy$-plane because of $h$.
@@ IV.2 Magnetic structures
1/3-plateau phase. In Figs. 9(a) and 9(b), we find a 1/3-plateau phase at around $\phi/\pi\sim 0.2$, which has a typical up-up-down configuration of spins. Figure 11(a) shows the wave number $q$ at which the Fourier transform of $\langle S^{z}_{j}\rangle$ shows a peak. In our analysis, we applied the SSD Fourier transform $\langle S^{z}_{q}\rangle=\sum_{j=1}^{N}f(r_{j})\langle S^{z}_{j}\rangle e^{-iqr_{j}}/\sum_{j=1}^{N}f(r_{j})$ Kawano and Hotta (2022), successfully detecting the plateau region yielding $q=2\pi/3$.
VC and ND phases. These phases are basically the same as $h=0$. It is also common to the case of $\Gamma=0$ except that $\langle Q^{xy}_{i}Q^{xy}_{j}\rangle\neq 0$. The VC phase at $\Gamma=0$ is a one-component TLLHikihara et al. (2010). The ND phase sustains at small $h$ and because of the finite spin gap observed at $h=0$, and indeed the magnetization is zero $\langle S^{z}_{i}\rangle=0$.
Fluctuating F phase. The fluctuating F phase has a ferromagnetic order with its moments pointing in the $z$ direction and the small fluctuation in the $xy$ plane. Its origin is essentially a $\Gamma$ term because it vanishes at $\Gamma=0$. We determined the boundary of this phase by $\langle S^{x}_{i}S^{x}_{i+N/2}\rangle\neq 0$ (see Fig. 11(b)).
Canted magnetic phases. The F1, F2, UD1, UD2, and UUDD phases at $h=0$ developed magnetic orders in the $xy$-plane. In a magnetic field, these moments cant in the $z$-directions.
@@ V Discussion and Summary
In the regime where the $\Gamma$-term dominates, the model exhibits strong magnetic anisotropy, resulting in ferromagnetic (F) and antiferromagnetic (AF-UD, UUDD) orderings with moments aligned along $(x,y,z)=(1,\pm 1,0)$ directions, namely, in-plane and perpendicular to the chain. The application of a magnetic field tilts these moments off the plane.
@@
We now explore the relevance of our findings to the ytterbium-based rare-earth magnet, YbCuS2. The Yb ion in this material is surrounded by S6 octahedraGulay and Olekseyuk (2005), leading to the formation of a Kramers $\Gamma_{6}$ doublet effectively carrying spin-1/2 due to the interplay between octahedral crystal fields of S-ions and strong spin-orbit coupling interactions Ohmagari et al. (2020). The crystal structure of YbCuS2 is orthorhombic with space group $Pnma$ Kaneshima (2021), where octahedra share edges along the $a$-direction and every two octahedra share edges along the $c$-direction, forming a one-dimensional zigzag chain of Yb ions. Experimental observations of the magnetic susceptibility $\chi(T)$ reveal antiferromagnetic interactions among the spins. At low temperatures, YbCuS2 undergoes a first-order transition at $T_{O}=0.95$ K, characterized by a pronounced divergence in the specific heatOhmagari et al. (2020). In the low-temperature phase, incommensurate magnetic structures are detected through 63/65Cu-nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) measurements on polycrystalline samples Hori et al. (2023). The nuclear spin-lattice relaxation rate $1/T_{1}$ of 63/65Cu-NQR exhibits a $T$-linear behavior at $T<0.5$ K, suggesting the presence of gapless excitations Hori et al. (2023). Additionally, the magnetic-field-temperature phase diagram obtained experimentally shows a transition of the low-temperature phase to an up-up-down (UUD) phase at around 5 TOhmagari et al. (2020).
@@
Finally, we emphasize the theoretically intriguing aspect of the ND phase in our phase diagram. In the standard nematic phase observed in spin ladders, the one-particle magnon excitation is gapped while the two-magnon excitation is gaplessHikihara et al. (2008); Läuchli et al. (2006). It typically occurs due to the localization of magnons caused by frustration-induced hopping cancellation. In such cases, the two-magnon bound state tends to condense earlier than the individual magnons, particularly near the saturation fieldZhitomirsky and Tsunetsugu (2010). In our model, the spin gap or a quasi-one-magnon gap in the non-conserved $S^{z}$ indeed occurs by the localization of spin-1/2 singlet pairs due to the frustration effect, breaking the lattice symmetry. However, the corresponding one-magnon instability does not compete with the quasi-two-magnon instability, as the corresponding $S^{z}$-oriented magnetic order never occurs in our phase diagram. This fact allows the unusual coexistence of the spin gap and the nonmagnetic gapless excitation in a wide range of parameters in the phase diagram. As we have discussed in §.II.2.1, the ND phase has ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ but breaks the translational symmetry. The former breaks at the magnetic phase transition, and the latter further breaks or recovers, depending on which of the magnetic phase to enter. This kind of phenomena may have similarity with the layers of discussions on the several symmetry-breakings reported in the Heisenberg spin ladder with the Dzyaloshinskii-Moriya interactionPenc et al. (2007).
@@ Appendix A Exact solutions
MG line and multicritical point. The $J_{2}/J_{1}=\Gamma_{2}/\Gamma_{1}=1/2$ is an MG line hosting singlet product state as an exact eigenstate, $|\Psi_{\rm MG}\rangle=\prod_{j=1}^{N/2}|s_{j}\rangle$ with $|s_{j}\rangle=(|\!\uparrow\downarrow\rangle-|\!\downarrow\uparrow\rangle)/\sqrt{2}$ on $[2j-1,2j]$ sites. The $\theta=0$ limit is the Majumdar-Ghosh modelMajumdar and Ghosh (1969a, b), and the upper endpoint with with $\Gamma_{\gamma}/J_{\gamma}=\sqrt{3}$ is the multicritical point $(\phi=0.1476\pi,\;\theta=\pi/3)$, with a ground state degeneracy of $(N+2)^{2}/4$ (even $N$), $(N+1)(N+3)/4$ (odd $N$) for an open boundary. The exact ground state can be obtained up to $N\sim 100$, with $\sim 3000$-fold degeneracy. The details of the method is given in Ref.[Saito and Hotta, 2024b].
@@ Appendix B Mean field approach around a multicritical point
At around the multicritical point in the phase diagram, the $J_{1}$-$J_{2}$-$\Gamma_{1}$-$\Gamma_{2}$ model described by Eq.(1) suffers a competition among several orders. The magnetic orders to be considered here are ferromagnet (F), two antiferromagnets with UD2 and UUDD configurations, and to accommodate all of them on equal footing, we need to consider the periodicity of magnetic moments up to four sites. The coordinate of the lattice sites are taken as $R=(4n-3)a,\cdots,4na$ with lattice constant $a$ where $n=1,\cdots,N/4$ is the unit cell index as shown in Fig. 12(a). We rely on the philosophy that the fluctuations around each order take place in a way that the magnetic moments belonging to the same magnetic sublattice are predominantly ferromagnetic.
@@
$E_{1x}=\frac{N}{4}(J_{1}-\Gamma_{1}+J_{2}-\Gamma_{2}),\quad\quad(\text{F})$
@@
We note that Eq.(25) provides the same energies as the energy expectation values of the original Hamiltonian Eq.(1) about the trial product wave functions of the corresponding F1 phase, the AF-UD2 phase the UUDD phase, e.g. $E_{1x}=\langle\Psi_{\rm F}|{\cal H}|\Psi_{\rm F}\rangle$
$|\Psi_{\rm F}\rangle=\prod_{j=1}^{N}|f\rangle_{j},\quad|\Psi_{\rm UD}\rangle=\prod_{j=1}^{N/2}|u\rangle_{2j-1}|d\rangle_{2j},$
@@ Appendix C Spin gap for various directions of magnetic field
Because the $z$-axis is a magnetic easy axis, the application of a magnetic field to the $z$-direction shown in Fig. 5 may naturally open a spin gap. Whereas, the field applied inside the $xy$ plane may not necessarily be the case because the system does not magnetically order inside the $xy$ plane in the ND phase. We thus apply a field in the four different directions, $\alpha=x,y,x^{\prime},$ and $y^{\prime}$, where the latter two are the magnetic easy axes of the F1,UD1 ($x^{\prime}$) and F2, UD2, UUDD($y^{\prime}$). We add the Zeeman term in these directions, $-h\sum_{i}S_{i}^{\alpha}$, and measure $M_{\alpha}$ using SSD-DMRG. Figure 13 shows the magnetization curves for the four cases. We find that ND phase has a robust spin gap which does not depend much on the field direction, confirming the finite spin gap. We also show that TLL and UUDD phases are gapless. In the F1 phase, the finite magnetization in the $x^{\prime}$-direction from $h=0$ is observed.
Figure 13: Magnetization curves for field $h$ applied along the (a) $x$, (b) $y$, (c) $x^{\prime}$, and (d) $y^{\prime}$-directions. We choose $(\phi,\theta)/\pi=(0.05,0.05)$ for TLL, $(0.2,0.05)$ for ND, $(0.15,0.4)$ for F1, and $(0.45,0.4)$ for AF-UUDD phases.

### v2

@@ III.4 Phase diagram
Figure 12: (a) Local magnetization $\langle S^{z}_{i}\rangle$ (upper panel) in the 1/3-plateau phase, clearly showing the up-up-down spin configuration. The peak wave number $q$ of ${\cal S}^{zz}(q)$ is shown as a function of $h$, where, within $0.5\pi\leq k\leq\pi$ we find $q=2\pi/3$. (a) The spin correlation function, $\langle S^{x}_{i}S^{x}_{i+N/2}\rangle\neq 0$ ($i=25$), as a function of $h$ that takes finite value in the fluctuating-FM phase at $1.5\lesssim h\lesssim 2.15$.
@@ IV Effect of magnetic field
It is known that the $J_{1}$-$J_{2}$ Heisenberg model in the AFM-AFM and FM-AFM regions exhibits a rich phase diagram in an applied magnetic fieldHikihara et al. (2008, 2010). Here, our $\Gamma$-terms change these phase diagrams a lot. We consider an external magnetic field along the $z$-axis whose Hamiltonian is given as ${\cal H}+{\cal H}_{f}$ using Eq.(1) with
@@ IV.1 Magnetic phase diagram
Figure 10 shows the magnetic phase diagram at fixed values of $\theta=0.05\pi,0.1\pi,0.3\pi,0.5\pi$ ($\Gamma/J\approx 0.16,0.32,1.4,\infty$). We first focus on small $\Gamma$, namely $\theta=0.05\pi,0.1\pi$. Compared to the zigzag Heisenberg model, the $\Gamma$-term transforms some of the nonmagnetic phases to magnetic ordered ones. For example, the fluctuating-FM phase at around $(\phi,\theta,h)\sim(0.2,0.05,1.8)$ ($J_{2}/J_{1}\approx 0.73,\Gamma/J\approx 0.16$) in Fig. 10(a) was originally a TLL2 phaseHikihara et al. (2010) but was stabilized as an emergent small magnetic order in the $xy$ plane. Furthermore, in the FM-AFM region $0.5\pi<\phi<\pi$, the canted UUDD phase where up-up-down-down magnetic order emerges in the $xy$ plane, was originally a vector chiral, nematic, and other multipolar phases when $\Gamma=0$.
At large $\Gamma$, i.e. $\theta=0.3\pi,0.5\pi$ in Figs. 10(c) and 10(d), the nonmagnetic phases like TLL and ND are almost wiped out. Magnetic phases, FM and UD, that appeared in the $h=0$ phase diagram are preserved while the moments are canted off the $xy$-plane because of $h$.
@@ IV.2 Magnetic structures
1/3-plateau phase. In Figs. 10(a) and 10(b), we find a 1/3-plateau phase at around $\phi/\pi\sim 0.2$ ($J_{2}/J_{1}\sim 0.73$), which has a typical up-up-down configuration of spins. Figure 12(a) shows the wave number $q$ at which the Fourier transform of $\langle S^{z}_{j}\rangle$ shows a peak. In our analysis, we applied the SSD Fourier transform $\langle S^{z}_{q}\rangle=\sum_{j=1}^{N}f(r_{j})\langle S^{z}_{j}\rangle e^{-iqr_{j}}/\sum_{j=1}^{N}f(r_{j})$ Kawano and Hotta (2022), successfully detecting the plateau region yielding $q=2\pi/3$.
VC and ND phases. These phases are basically the same as $h=0$. It is also common to the case of $\Gamma=0$ except that $\langle Q^{xy}_{i}Q^{xy}_{j}\rangle\neq 0$. The VC phase at $\Gamma=0$ is a one-component TLLHikihara et al. (2010). The ND phase sustains up to finite $h$ whose value corresponds to the spin gap at $h=0$.
Fluctuating FM phase. The fluctuating FM phase has a ferromagnetic order with its moments pointing in the $z$ direction and the small fluctuation in the $xy$ plane. Its origin is essentially a $\Gamma$ term because it vanishes at $\Gamma=0$. We determined the boundary of this phase by $\langle S^{x}_{i}S^{x}_{i+N/2}\rangle\neq 0$ (see Fig. 12(b)).
Canted magnetic phases. The FM1, FM2, UD1, UD2, and UUDD phases at $h=0$ developed magnetic orders in the $xy$-plane. In a magnetic field, these moments cant in the $z$-directions.
@@ V Discussion and Summary
In the regime where the $\Gamma$-term dominates, the model exhibits strong magnetic anisotropy, resulting in ferromagnetic (FM) and antiferromagnetic (AFM-UD, UUDD) orderings with moments aligned along $(x,y,z)=(1,\pm 1,0)$ directions, namely, in-plane and perpendicular to the chain. The application of a magnetic field tilts these moments off the plane.
@@
We now explore the relevance of our findings to the ytterbium-based rare-earth magnet, YbCuS2. The Yb ion in this material is surrounded by S6 octahedraGulay and Olekseyuk (2005), leading to the formation of a Kramers $\Gamma_{6}$ doublet effectively carrying spin-1/2 due to the interplay between octahedral crystal fields of S-ions and strong spin-orbit coupling interactions Ohmagari et al. (2020). The crystal structure of YbCuS2 is orthorhombic with space group $Pnma$ Kaneshima (2021), where octahedra share edges along the $a$-direction and every two octahedra share edges along the $c$-direction, forming a one-dimensional zigzag chain of Yb ions, whose details are shown in Ref.[Saito et al., 2024]. Experimental observations of the magnetic susceptibility $\chi(T)$ reveal antiferromagnetic interactions among the spins. At low temperatures, YbCuS2 undergoes a first-order transition at $T_{O}=0.95$ K, characterized by a pronounced divergence in the specific heatOhmagari et al. (2020). In the low-temperature phase, incommensurate magnetic structures are detected through 63/65Cu-nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) measurements on polycrystalline samples Hori et al. (2023). The nuclear spin-lattice relaxation rate $1/T_{1}$ of 63/65Cu-NQR exhibits a $T$-linear behavior at $T<0.5$ K, suggesting the presence of gapless excitations Hori et al. (2023). Additionally, the magnetic-field-temperature phase diagram obtained experimentally shows a transition of the low-temperature phase to an up-up-down (UUD) phase at around 5 TOhmagari et al. (2020).
@@
Finally, we emphasize the theoretically intriguing aspect of the ND phase in our phase diagram. In the standard nematic phase observed in spin ladders, the one-particle magnon excitation is gapped while the two-magnon excitation is gaplessHikihara et al. (2008); Läuchli et al. (2006). It typically occurs due to the localization of magnons caused by frustration-induced hopping cancellation. In such cases, the two-magnon bound state tends to condense earlier than the individual magnons, particularly near the saturation fieldZhitomirsky and Tsunetsugu (2010). In our model, the spin gap or a quasi-one-magnon gap in the non-conserved $S^{z}$ indeed occurs by the localization of spin-1/2 singlet pairs due to the frustration effect, breaking the lattice symmetry. However, the corresponding one-magnon instability does not compete with the quasi-two-magnon instability, as the corresponding $S^{z}$-oriented magnetic order never occurs in our phase diagram. This fact allows the unusual coexistence of the spin gap and the nonmagnetic gapless excitation in a wide range of parameters in the phase diagram. As we have discussed in §.II.2.2, the ND phase has ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ but breaks the translational symmetry. The former breaks at the magnetic phase transition, and the latter further breaks or recovers, depending on which of the magnetic phases to enter. This kind of phenomena may have similarity with the layers of discussions on the several symmetry-breakings reported in the Heisenberg spin ladder with the Dzyaloshinskii-Moriya interactionPenc et al. (2007).
@@ Appendix A Exact solutions
Majumdar-Ghosh(MG) line and multicritical point. The $J_{2}/J_{1}=\Gamma_{2}/\Gamma_{1}=1/2$ is a MG line hosting singlet product state as an exact eigenstate, $|\Psi_{\rm MG}\rangle=\prod_{j=1}^{N/2}|s_{j}\rangle$ with $|s_{j}\rangle=(|\!\uparrow\downarrow\rangle-|\!\downarrow\uparrow\rangle)/\sqrt{2}$ on $[2j-1,2j]$ sites. The $\theta=0$ limit is the Majumdar-Ghosh modelMajumdar and Ghosh (1969a, b), and the upper endpoint with with $\Gamma_{\gamma}/J_{\gamma}=\sqrt{3}$ is the multicritical point $(\phi=0.1476\pi,\;\theta=\pi/3)$, with a ground state degeneracy of $(N+2)^{2}/4$ (even $N$), $(N+1)(N+3)/4$ (odd $N$) for an open boundary. The exact ground state can be obtained up to $N\sim 100$, with $\sim 3000$-fold degeneracy. The details of the method are given in Ref.[Saito and Hotta, 2024b].
@@ Appendix B Mean field approach around a multicritical point
At around the multicritical point in the phase diagram, the $J_{1}$-$J_{2}$-$\Gamma_{1}$-$\Gamma_{2}$ model described by Eq.(1) suffers a competition among several orders. The magnetic orders to be considered here are ferromagnet (FM), two antiferromagnets with UD2 and UUDD configurations, and to accommodate all of them on equal footing, we need to consider the periodicity of magnetic moments up to four sites. The coordinate of the lattice sites are taken as $R=(4n-3)a,\cdots,4na$ with lattice constant $a$ where $n=1,\cdots,N/4$ is the unit cell index as shown in Fig. 13(a). We rely on the philosophy that the fluctuations around each order take place in a way that the magnetic moments belonging to the same magnetic sublattice are predominantly ferromagnetic.
@@
$E_{1x}=\frac{N}{4}(J_{1}-\Gamma_{1}+J_{2}-\Gamma_{2}),\quad\quad(\text{FM})$
@@
We note that Eq.(32) provides the same energies as the energy expectation values of the original Hamiltonian Eq.(1) about the trial product wave functions of the corresponding FM1 phase, the AFM-UD2 phase the UUDD phase, e.g. $E_{1x}=\langle\Psi_{\rm FM}|{\cal H}|\Psi_{\rm FM}\rangle$
$|\Psi_{\rm FM}\rangle=\prod_{j=1}^{N}|f\rangle_{j},\quad|\Psi_{\rm UD}\rangle=\prod_{j=1}^{N/2}|u\rangle_{2j-1}|d\rangle_{2j},$
@@ Appendix C Spin gap for various directions of magnetic field
Because the $z$-axis is a magnetic easy axis, the application of a magnetic field to the $z$-direction shown in Fig. 5 may naturally open a spin gap. Whereas, the field applied inside the $xy$ plane may not necessarily be the case because the system does not magnetically order inside the $xy$ plane in the ND phase. We thus apply a field in the four different directions, $\alpha=x,y,x^{\prime},$ and $y^{\prime}$, where the latter two are the magnetic easy axes of the FM1,UD1 ($x^{\prime}$) and FM2, UD2, UUDD($y^{\prime}$). We add the Zeeman term in these directions, $-h\sum_{i}S_{i}^{\alpha}$, and measure $M_{\alpha}$ using SSD-DMRG. Figure 14 shows the magnetization curves for the four cases. We find that ND phase has a robust spin gap which does not depend much on the field direction, confirming the finite spin gap. We also show that TLL and UUDD phases are gapless. In the FM1 phase, the finite magnetization in the $x^{\prime}$-direction from $h=0$ is observed.
Figure 14: Magnetization curves for field $h$ applied along the (a) $x$, (b) $y$, (c) $x^{\prime}$, and (d) $y^{\prime}$-directions. We choose $(\phi,\theta)/\pi=(0.05,0.05)$ for TLL, $(0.2,0.05)$ for ND, $(0.15,0.4)$ for FM1, and $(0.45,0.4)$ for AFM-UUDD phases.